Monthly payments of $50 were deposited at the end of each month for five years. The account earned 6.6% compounded monthly. Determine the balance three years after the last deposit.
1
Expert's answer
2020-11-09T20:35:23-0500


find monthly rate:

"\frac{6.6}{100}=0.066"

"\frac{0.055}{12}=0.0055"

1.055=(1+1.0055)^n -

increment factor

substitute the monthly rate into the postnumerando annuity formula

this is a postnumerando annuity:

"FV=50(\frac{1.0055^{36}-1}{0.0055})=165.38"